# Kairosis: A method for dynamical probability forecast aggregation informed by Bayesian change point detection

### v2

@@ 3.2 Performance Evaluation
In each case (except (5), where they are equal) we compute both the weighted mean and the weighted median (that is, for rank-ordered forecasts $i=1,\ldots,N$ with normalized weights $w_{i}$, the forecast $f_{j}$ with $j$ the smallest integer such that $\sum_{i=1}^{j}w_{i}>{1\over 2}$).
@@ Appendix A A worked example of computing the kairosis weighting function
$\underbrace{P(t_{\rm\small CP}=t\mid\text{Forecasts})}_{\text{Posterior}}\propto\underbrace{P(\text{Forecasts}\mid t_{\rm\small CP}=t)}_{\text{likelihood}}\underbrace{P(t_{\rm\small CP}=t)}_{\text{Prior}}$ $t\in\{t_{1},t_{2},t_{3}\}.$
We specify a prior probability that decays geometrically as explained in Section 2.1. The likelihood term factorizes into two parts on the basis that a change point is assumed to force a change in the distribution of the forecasts before and after it so that
$P(\text{Forecasts}\mid t_{\rm\small CP}=t)=$ $P(\text{Forecasts made before }t\mid t_{\rm\small CP}=t)$
×$\times P(\text{Forecasts made after }t\mid t_{\rm\small CP}=t).$
The factors here are based on the bin counts before and after the proposed change point $t$. They are computed using the probability mass function of the Dirichlet-categorical distribution whose functional form is provided in equation (4). This mass function is maximized when all the counts fall within a single bin and minimized when they are evenly distributed across them all. Intermediate values are attained for intermediate degrees of forecast dispersion.
@@ Appendix B Non-Probabilistic Questions
The crucial alteration of the method is in the binning of forecasts. Probability forecasts are all on the same, fixed interval $(0,1)$, whereas with general point forecasts the domain varies from question to question, necessitating bins which vary not only between questions but also with present (forecaster) time $N$. In our experimental calculations, whose results are presented below, we adopted a simple adaptive binning strategy which involved partitioning the range of possible forecast values according to the quantiles of the forecasts observed up until the aggregated forecast is needed. This partition defines the bin structure that is used to compare forecast distributions to either side of each candidate change point which, following the calculations described in the main paper, leads the our kairosis weighting function used for aggregation.
@@ Appendix C Sensitivity Analysis
We now assess the sensitivity of our kairosis calculations to changes in the parameters $p$, which governs the geometric prior on change point locations, and $\lambda$, which governs the rate at which pseudo-counts $\mathbb{\alpha}_{k}=\lambda t_{\rm\small CP}$ are added to the Dirichlet prior for bin probabilities before the most recent change point. Specifically, we calculate Brier scores for aggregated forecasts using weights derived from our kairosis calculations and average these over the Metaculus questions. The calculations are repeated for a range of values of $\{p,\lambda\}$ and the results plotted in Figure 7.

### v4

@@ 3.2 Performance Evaluation
In each case (except (5), where they are equal) we compute both the weighted mean and the weighted median (that is, for forecasts $i=1,\ldots,R$ ordered by increasing weights $w_{i}$, the forecast $f_{j}$ with $j$ the smallest integer such that $\sum_{i=1}^{j}w_{i}>{1\over 2}$).
@@ Appendix A A worked example of computing the kairosis weighting function
$\underbrace{P(t_{\rm\small CP}=t_{r}\mid\text{Forecasts})}_{\text{posterior}}\propto\underbrace{P(\text{Forecasts}\mid t_{\rm\small CP}=t_{r})}_{\text{likelihood}}\underbrace{P(t_{\rm\small CP}=t_{r})}_{\text{prior}}$ $t_{r}\in\{t_{1},t_{2},t_{3}\}.$
We specify a prior probability that decays geometrically as explained in Section 2.1. The likelihood term factorizes into two parts on the basis that a change point is assumed to force a change in the distribution of the forecasts up to and after it so that
$P(\text{Forecasts}\mid t_{\rm\small CP}=t_{r})=$ $P(\text{Forecasts made up to }t\mid t_{\rm\small CP}=t_{r})$
×$\times P(\text{Forecasts made after }t\mid t_{\rm\small CP}=t_{r}).$
The factors here are based on the bin counts up to and after the proposed change point $t_{r}$. They are computed using the probability mass function of the Dirichlet-categorical distribution whose functional form is provided in equation (4). This mass function is maximized when all the counts fall within a single bin and minimized when they are evenly distributed across them all. Intermediate values are attained for intermediate degrees of forecast dispersion.
@@ Appendix B Non-Probabilistic Questions
The crucial alteration of the method is in the binning of forecasts. Probability forecasts are all on the same, fixed interval $[0,1]$, whereas with general point forecasts the domain varies from question to question, necessitating bins which vary not only between questions but also with present (forecaster) time $R$. In our experimental calculations, whose results are presented below, we adopted a simple adaptive binning strategy which involved partitioning the range of possible forecast values according to the quantiles of the forecasts observed up until the aggregated forecast is needed. This partition defines the bin structure that is used to compare forecast distributions to either side of each candidate change point which, following the calculations described in the main paper, leads to our kairosis weighting function used for aggregation.
@@ Appendix C Sensitivity Analysis
We now assess the sensitivity of our kairosis calculations to changes in the parameters $p$, which governs the geometric prior on change point locations, and $\lambda$, which governs the rate at which pseudo-counts $\mathbb{\alpha}_{k}=\lambda t_{\rm\small CP}$ are added to the Dirichlet prior for bin probabilities before the supposed most recent change point. Specifically, we calculate Brier scores for aggregated forecasts using weights derived from our kairosis calculations and average these over the Metaculus questions. The calculations are repeated for a range of values of $\{p,\lambda\}$ and the results plotted in Figure 7.